OpenStudy (anonymous):

MEDAL! Please explain clearly, step by step how to do this! Solve for x. 5a/x = 5b/x-1

pooja195 (pooja195):

\[\huge~\rm~\frac{ 5a }{ x}=\frac{ 5b }{ x-1}\]

pooja195 (pooja195):

like that?

OpenStudy (anonymous):

@pooja195 Yeah! :)

pooja195 (pooja195):

Do you know where to start? :)

OpenStudy (anonymous):

Don't I cross multiply?

pooja195 (pooja195):

yes! :) set uo the equation u get once you do that

OpenStudy (anonymous):

5a(x-1) = 5b(x) Is this right so far? Or am I way off...?

pooja195 (pooja195):

good continue

OpenStudy (anonymous):

Whats the next step?

pooja195 (pooja195):

5a(x-1) distribute

OpenStudy (anonymous):

How do I do that? I'm sorry I just suck at math... Every other subject are A's! XD

pooja195 (pooja195):

\[\huge~\rm~5a~\times~x=?\] \[\huge~\rm~5\times-1=? \]

OpenStudy (anonymous):

5ax and 4x?

pooja195 (pooja195):

we are multiplying so it would be -5 5ax−5a=5bx

pooja195 (pooja195):

what do u do from here?

pooja195 (pooja195):

need help ?

OpenStudy (anonymous):

Yes please

pooja195 (pooja195):

hmm :/ im not sure where to go from here maybe @mLe can help ;-; ?

OpenStudy (mle):

get all the terms that have x in it alone to one side

OpenStudy (mle):

subtract 5bx from both sides

OpenStudy (mle):

and add 5a to both sides

OpenStudy (mle):

then move the x out to the front ... opposite of distributing... for example if I had 7xy + 9xz I would get x (7y + 9z) then divide both sides by the stuff in the parenthesis to get it alone